OpenStudy (anonymous):

(-3,1) y = -1/3x What are the steps to determine if the ordered pair is a solution of the equation? More specifically, How do I multiply -1/3 by -3. The negatives are confusing me.

OpenStudy (phi):

See http://www.khanacademy.org/math/arithmetic/absolute-value/mult_div_negatives/v/why-a-negative-times-a-negative-is-a-positive a minus times a minus is a positive to multiply the fractions (think of -3 as -3/1) you multiply top times top and bottom times bottom \[ -\frac{3}{1} \cdot - \frac{1}{3} = + \frac{3\cdot 1}{1\cdot 3}= \frac{3}{3}\] 3/3 is 1

OpenStudy (phi):

do you know how to multiply 3 * 1/3 ?